1. If the height of adults in Illinois State University can be modelled using H~ N(174,35.7). Where the height is measured in centimeters

a. What type of data is this? (4 points)

b. What is the probability that a randomly selected adult from the University has height between 166 and 183? (14 points, show your work)


1
Expert's answer
2022-03-18T00:55:33-0400

Solution

(a) This is a normal distributed continuous data since the height data falls in a constant sequence.


(b) Probability that "166<Z<183"

Population mean "\mu =174"

Standard deviation "\sigma =35.7"


"Z=\dfrac{X-\mu}{\sigma}"


"P(\frac{166-174}{35.7}<Z<\frac{183-174}{35.7})"


"P(-0.224<Z<0.252)"


"=P(Z<0.252)-P(Z<-1.5)"


"=0.5995-0.4114"


"=0.1881"